A step-down transformer produces a voltage of 5.9 V across the secondary coil when the voltage across the primary coil is 140 V .

Part A What voltage appears across the primary coil of this transformer if 140 V is applied to the secondary coil?

Express your answer using two significant figures.
Answer #1

The primary to secondary turns ratio in an ideal transformer is equal to the ratio of primary voltage to the secondary voltage

Vp/Vs = Np/Ns

140/5.9 = Np/Ns

Vp/140 = Np/Ns

Vp = 140*140/5.9

Vp = 3322.0339 V